A recipe for rice pudding uses 1/2 cup of rice and serves 6 people. How many cups of rice do you need to serve 75 people?
Answer by stanbon(75887)

A recipe for rice pudding uses 1/2 cup of rice and serves 6 people. How many cups of rice do you need to serve 75 people?
-------
Proportion:
rice/people = rice/people
----------------------
x/75 = (1/2)/6
---
x = 75(1/2)/6
---
x = 75/12
---
x = 6 1/4 cups
======================
Cheers,
Stan H.